OpenStudy (anonymous):

Photocopying. Quick Copy recently charged 49¢ per page for color copies and 7¢ per page for black-and-white copies. If Shirlee’s bill for 90 copies was $11.34, how many copies of each type were made?

OpenStudy (whpalmer4):

Let C be number of color copies and B be the number of black and white copies. Total copies is 90, so B+C=90. 0.49C + 0.07B = 11.34 to tally the costs. Solve for B and C

OpenStudy (texaschic101):

c = colored copies b = black and white copies c + b = 90 -->c = 90 - b .49c + .07b = 11.34 now just sub 90 - b in for c in the 2nd equation .49(90 - b) + .07b = 11.34 can you take it from here and solve for b ? And once b is found, sub that back into the 1st equation and solve for c.

OpenStudy (whpalmer4):

I would multiply second equation by 100 to eliminate decimals, and solve first equation to give one variable in terms of the other, then substitute that in second equation.